OpenStudy (tiffany_rhodes):

First make a substitution and then using IBP to evaluate the integral -3sin(x^1/2). I was going to use x^(1/2) as U and let du= 2/3x^(3/2) dx. I'm not sure how to incorporate u-sub. with IBP though.

OpenStudy (anonymous):

What is the integral you are trying to solve?

OpenStudy (anonymous):

\[-3 \int\limits \sin(x^{1/2})dx?\]

OpenStudy (anonymous):

The substitution isn't as straight-forward as you may be used to. \[-3\int\sin\sqrt x~dx\] Let \(\sqrt x=u\), so that \(x=u^2\). This gives you \(dx=2u~du\). \[-3\int\sin u~(2u~du)\\ -6\int u\sin u~du\]

OpenStudy (tiffany_rhodes):

Thank you! I'll see where I get from here.

OpenStudy (anonymous):

You're welcome